OpenStudy (anonymous):

Hi, could someone tell me how they got an answer of: x/12 for the question below: A boat costs x dollars, and this cost is to be shared equally by a group of people. In terms of x, how many dollars less will each person contribute if there are 4 people in the group instead of 3.

OpenStudy (anonymous):

If there are 3 people, each will contribute (x/3) $ If there are 4, each will contribute (x/4) $ The question is how much less the individual contribution will be if there were 4 peopl instead of 3 = (x/3) - (x/4) = (x/12)

OpenStudy (anonymous):

but why would it equal to (x/12)?

OpenStudy (anonymous):

take LCM x/3-x/4 LCM is 12 \[\Large \frac{x}{3}-\frac{x}{4}=\frac{4x-3x}{12}=\frac{x}{12}\]

OpenStudy (anonymous):

oh well that makes a lot more sense c: thank you

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

If there are 3 people,each will pay $ x/3 If there are 4 people,each will pay $ x/4 your question says that 'how many dollars less will each person contribute if there are 4 people in the group instead of 3'.This means we should subtract x/4 from x/3 \[\frac{ x }{ 3 }-\frac{ x }{ 4}\],take lcm,lcm=12 x/3-x/4=4x-3x/12=x/12

OpenStudy (anonymous):

tell carl to go away he made my friend spaghetti jr cry